find all solutions of the system.
x+2y=4
2x+4y=8

Answer by funmath(2933)

find all solutions of the system.
x+2y=4
2x+4y=8
This has an infinite number of solutions, notice that if you multiply the top equation by 2 you'll get the bottom equation. These two line lie right on top of each other.
If you didn't notice that, here's what happens when you try to eliminate the x's by multiplying the top equation by -2 and adding the two equations together.
-2(x+2y=4) ----2x-4y=-8
:
-2x-4y=-8
2x+4y=8
----------
0+0=0
0=0 This tells you that there are an infinite number of solutions.
Happy Calculating!!!!